5.8 You are considering starting a walk in clinic. Your financial projections for the first year of operations are as follows: Revenues (10,000) $400,000 Wages and Benefits $220,000 Rent Depreciation 5,000 30,000 2,500 50,000 Medical Supplies Administrative Supplies 10,000 Assume that all cost are fixed except supply costs, which are variable. Furthermore, assume that the linic must pay taxes at a 30 percent rate Construct the clinics projected P & L statement. What number of visit is required to break-even What number of visits is re This problem has been solved! Expert Answer G See the answer acer Cl
Answer #1

Answer to Requirement 1 P&L Statement Revenue Expenses: Wages and benefits Rent Depreciation Utilities Medical Supplies Administrative Supplies Total Expenses Taxable Income Taxes Net Income 400000 220000 5000 30000 2500 50000 10000 317500 82500 24750 57750

Answer to Requirement 2.

Break Even Point (Number of Visits) = Fixed Cost / Contribution Margin per Unit

Total Variable Cost = $50,000 + $10,000 = $60,000
Variable Cost per Unit = $60,000 / 10,000 = $6
Sales per Unit = $400,000 / 10,000 = $40
Contribution Margin per Unit = $40 - $6 = $34

Fixed Cost = $220,000 + $5,000 + $30,000 + $2,500
Fixed Cost = $257,500

Break Even Point (Number of Visits) = 257,500 / 34
Break Even Point (Number of Visits) = 7,574 visits

Requirement 3. could not be done, as complete question is not visible.